# Solutions for $k$-generalized Fibonacci numbers using Fuss-Catalan numbers

### v3

@@ Abstract
We present new expressions for the $k$-generalized Fibonacci numbers, say $F_{k}(n)$. They satisfy the recurrence $F_{k}(n)=F_{k}(n-1)+\dots+F_{k}(n-k)$. Explicit expressions for the roots of the auxiliary (or characteristic) polynomial are presented, using Fuss-Catalan numbers. Properties of the roots are enumerated. We quantify the accuracy of asymptotic approximations for $F_{k}(n)$ for $n\gg 1$. Our results subsume and extend some results published by previous authors. We also comment on the use of multinomial sums for the $k$-generalized Fibonacci numbers and related sequences. We also employ the generating function to express $F_{k}(n)$ as a concise (non-nested) sum of binomial coefficients for arbitrary $k$. Finally, we present a basis (or ‘fundamental solutions’) to solve the above recurrence for arbitrary initial conditions.
@@ 1 Introduction
We treat the so-called ‘$k$-generalized’ Fibonacci numbers, say $F_{k}(n)$. They satisfy the recurrence
@@
The usual Fibonacci numbers are the special case $k=2$. Eq. (1.1) is a homogeneous linear difference equation and its characteristic equation is
@@
The corresponding auxiliary (or characteristic) polynomial is $A(x)=x^{k}-x^{k-1}-\dots-1$. Several authors [1, 2, 3] have shown that the roots of $A(x)$ are simple. Let us denote the roots of $A(x)$ by $\zeta_{\ell}$, $\ell=0,\dots,k-1$. Then we express $F_{k}(n)$ as a weighted sum of powers of the roots as follows (i.e. a Binet-style formula)
$F_{k}(n)=\sum_{\ell=0}^{k-1}c_{\ell}\zeta_{\ell}^{n}\,.$ (1.3)
Here $\{c_{0},\dots,c_{k-1}\}$ is a set of coefficients which do not depend on $n$, but they do depend on $k$ and the initial conditions, i.e. the values of $F_{k}(0)$ through $F_{k}(k-1)$. It is standard to employ the initial conditions $F_{k}(0)=\dots=F_{k}(k-2)=0$ and $F_{k}(k-1)=1$. We shall begin with this, but later we shall treat alternative initial conditions, by Spickerman and Joyner [4]. In Sec. 8, we present a complete set of ‘fundamental solutions’ for the recurrence, i.e. a basis which can be used to solve for arbitrary initial conditions.
Our main result is to present explicit expressions for the roots $\zeta_{j}$, in terms of so-called ‘Fuss-Catalan numbers’ (see the text by Graham, Knuth and Patashnik [5] for relevant definitions, formulas and identities). We also employ the generating function to express $F_{k}(n)$ as a concise (non-nested) sum of binomial coefficients for arbitrary $k$ (see Sec. 7). Finally, in Sec. 8 we present a basis (or ‘fundamental solutions’) to solve the above recurrence for arbitrary initial conditions. The connection to Fuss-Catalan numbers is somewhat surprising. Dilworth and Mane [6] (hereafter ‘DM’) published a paper on success run waiting times of Bernoulli trials. The recurrence for the probability mass function $f_{k}(n)$ is ([6] eq. (1)), with $p\in(0,1)$, $q=1-p$ and $k\geq 2$,
@@
Set $p=q=\frac{1}{2}$ and $f_{k}(n)=F_{k}(n)/2^{n}$; then elementary manipulations yield eq. (1.1). Hence the solution for the $k$-generalized Fibonacci numbers $F_{k}(n)$ equals $2^{n}$ times the probability mass function $f_{k}(n)$ of success run waiting times of Bernoulli trials, with $p=q=\frac{1}{2}$. The auxiliary polynomial associated with eq. (1.4) is $\mathcal{A}_{p}(x)=x^{k}-qx^{k-1}-pqx^{k-2}-\cdots-p^{k-1}q$. Barry and Lo Bello [7] showed that the roots of $\mathcal{A}_{p}(x)$ are simple, which goes to show that there are fruitful connections between different fields of research. Let us denote the roots of $\mathcal{A}_{p}(x)$ by $\lambda_{\ell}$, $\ell=0,\dots,k-1$. Then $\zeta_{\ell}=2\lambda_{\ell}$, with $p=q=\frac{1}{2}$.
It is known that $\mathcal{A}_{p}(x)$ has a unique positive root, which has a larger magnitude than all the other roots [8]. We denote it by $\lambda_{0}$ below. It is also known that if $k$ is even then there is a unique negative root. For all $k$, the remaining roots come in complex conjugate pairs [6]. Wolfram [3] also proved these results for the case of $k$-generalized Fibonacci numbers. DM employed the terms “primary root” for $\lambda_{0}$ and “secondary roots” for the other $\lambda_{j}$, $j=1,\dots,k-1$.
@@
The primary root $\zeta_{0}$ is given by ([6] eq. (8))
@@
This agrees with eq. (1.8); the proof is as follows.
@@
It remains now to derive expressions for $F_{k}(n)$ for specific initial conditions, and also to derive asymptotic expressions for $n\gg 1$ (in which case the solution is dominated by the contribution from the principal root). To do so, it will be helpful below to work with the polynomial $B(x)=(x-1)A(x)$ for the following reason. Note that
@@
Remark 1.1.
Gabai ([9] Theorem 1) published the following recurrence, which can also be deduced from $B(x)$ above
$F_{k}(n+1)=2F_{k}(n)-F_{k}(n-k)\,.$ (1.14)
For all $k\geq 2$, there are only two terms on the right-hand side. Initial values must be supplied for $n=0,\dots,k$. Eq. (1.14) has the extraneous solution $F_{k}(n)=1$ for all $n$; the choice of initial values will exclude this case.
@@ 2 Properties of roots
Distribution of the roots in the complex plane. Observe the following limits as $k\to\infty$: (i) for the principal root, $\zeta_{0}\to 2^{-}$ (from eq. (2.3)) and (ii) for all the secondary roots, $|\zeta_{j}|\to 1^{-}$ (from eq. (2.4)). Hence all the secondary roots approach the unit circle (from the inside), but the principal root approaches $2$ (from below). A graph of the roots in the complex plane is plotted in Fig. 1, for $k=40$. For $k\gg 1$, the secondary roots are also distributed approximately uniformly around the unit circle. A graph of the arguments of the secondary roots, i.e. $\arg(\zeta_{j})/(2\pi)$, is plotted in Fig. 2, for $k=40$. The points lie close to a straight line.
@@ 3 Vandermonde
First recall eq. (1.3). We set up a Vandermonde system of equations to solve for the coefficients $c_{\ell}$ as follows
$\begin{split}\begin{pmatrix}1&1&\dots&1\\ \zeta_{0}&\zeta_{1}&\dots&\zeta_{k-1}\\ \zeta_{0}^{2}&\zeta_{1}^{2}&\dots&\zeta_{k-1}^{2}\\ \vdots&&&\vdots\\ \zeta_{0}^{k-1}&\zeta_{1}^{k-1}&\dots&\zeta_{k-1}^{k-1}\end{pmatrix}\begin{pmatrix}c_{0}\\ c_{1}\\ c_{2}\\ \vdots\\ c_{k-1}\end{pmatrix}=\begin{pmatrix}F_{k}(0)\\ F_{k}(1)\\ F_{k}(2)\\ \vdots\\ F_{k}(k-1)\end{pmatrix}\,.\end{split}$ (3.1)
We begin with the initial conditions $F_{k}(0)=\dots=F_{k}(k-2)=0$ and $F_{k}(k-1)=1$. (We treat the initial condition by Spickerman and Joyner [4] in Sec. 5.) The solution is
$c_{\ell}=\frac{1}{\prod_{j\neq\ell}(\zeta_{\ell}-\zeta_{j})}\,.$ (3.2)
@@
$B^{\prime}(x)=\frac{B(x)}{x-1}+\sum_{j=0}^{k-1}\frac{B(x)}{x-\zeta_{j}}\,.$ (3.4)
Then, noting that $B(\zeta_{\ell})=0$,
$B^{\prime}(\zeta_{\ell})=(\zeta_{\ell}-1)\prod_{j\neq\ell}(\zeta_{\ell}-\zeta_{j})\,.$ (3.5)
Also $B^{\prime}(x)=(k+1)x^{k}-2kx^{k-1}$, thus $B^{\prime}(\zeta_{\ell})=((k+1)\zeta_{\ell}-2k)\zeta_{\ell}^{k-1}$. Hence
$\begin{split}c_{\ell}&=\frac{1}{\prod_{j\neq\ell}(\zeta_{\ell}-\zeta_{j})}=\frac{\zeta_{\ell}-1}{B^{\prime}(\zeta^{\ell})}=\frac{\zeta_{\ell}-1}{((k+1)\zeta_{\ell}-2k)\zeta_{\ell}^{k-1}}\,.\end{split}$ (3.6)
@@
$F_{k}(n)=\sum_{\ell=0}^{k-1}\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\zeta_{\ell}^{n-k+1}\,.$ (3.7)
@@ 5 Alternative initial conditions: Spickerman and Joyner
Spickerman and Joyner [4] solved the recurrence eq. (1.1) as follows. First to avoid confusions of notation, we follow them and write $u_{n}$ in place of $F_{k}(n)$. They employed the initial conditions $(u_{0},\dots,u_{k-1})=(1,1,2,4,\dots,2^{k-2})$, i.e. $u_{0}=1$ and $u_{n}=2^{n-1}$ for $11\leq n\leq k-$ 1. Hence the next term is $u_{k}=1+1+2+4+\dots+2^{k-2}=2^{k-1}$. Let us change this slightly and shift the indices down by one (and discard $u_{0}$) and say instead $(v_{0},\dots,v_{k-1})=(1,2,4,\dots,2^{k-1})$, i.e. $v_{n}=2^{n}$ for $10\leq n\leq k-$ 1. Let us generalize this even more and replace $2$ by $\mu$, and write $w_{n}=\mu^{n}$ for $10\leq n\leq k-$ 1. Then the solution for $c_{\ell}$ in the Vandermonde matrix equations in Sec. 3 is easy; it is the Lagrange polynomial $c_{\ell}=L_{\ell}(\mu)$. The Lagrange polynomial $L_{\ell}(x)$ is
$L_{\ell}(x)=\frac{(x-\zeta_{0})\dots(x-\zeta_{\ell-1})(x-\zeta_{\ell+1})\dots(x-\zeta_{k-1})}{(\zeta_{\ell}-\zeta_{0})\dots(\zeta_{\ell}-\zeta_{\ell-1})(\zeta_{\ell}-\zeta_{\ell+1})\dots(\zeta_{\ell}-\zeta_{k-1})}\,.$ (5.1)
This can be obtained from the auxiliary polynomial, written in the form (see eq. (1.13)) $A(x)=(x^{k+1}-2x^{k}+1)/(x-1)$. The denominator of $L_{\ell}(x)$ is $D_{\ell}=A^{\prime}(\zeta_{\ell})$ and does not depend on $x$. The numerator of $L_{\ell}(\mu)$ is $N_{\ell}(\mu)=A(\mu)/(\mu-\zeta_{\ell})$, provided $\mu$ is not a root of $A(x)$. For now, we assume $\mu$ is not a root of $A(x)$. Then the numerator is
$N_{\ell}(\mu)=\frac{\mu^{k+1}-2\mu^{k}+1}{(\mu-1)(\mu-\zeta_{\ell})}\,.$ (5.2)
@@
The denominator is given by (note that the second term in $A^{\prime}(x)$ vanishes for $x=\zeta_{\ell}$)
$D_{\ell}=\frac{(k+1)\zeta_{\ell}^{k}-2k\zeta_{\ell}^{k-1}}{\zeta_{\ell}-1}=\frac{(k+1)\zeta_{\ell}-2k}{\zeta_{\ell}-1}\zeta_{\ell}^{k-1}\,.$ (5.4)
Hence the coefficient $c_{\ell}$ is
$c_{\ell}=\frac{N_{\ell}(\mu)}{D_{\ell}}=\frac{\mu^{k+1}-2\mu^{k}+1}{(\mu-1)(\mu-\zeta_{\ell})}\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\frac{1}{\zeta_{\ell}^{k-1}}\,.$ (5.5)
@@
$\lim_{\mu\to 1}c_{\ell}=\frac{1-k}{(1-\zeta_{\ell})}\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\frac{1}{\zeta_{\ell}^{k-1}}=\frac{k-1}{(k+1)\zeta_{\ell}-2k}\frac{1}{\zeta_{\ell}^{k-1}}\,.$ (5.6)
@@
What if $\mu$ is a root of $A(x)$? Actually this is easy. If $\mu=\zeta_{i}$ then by construction of the Lagrange polynomials, $L_{i}(\zeta_{i})=1$ and $L_{\ell}(\zeta_{i})=0$ for $i\neq\ell$. Hence $c_{i}=1$ and $c_{\ell}=0$ for $i\neq\ell$, i.e. $c_{\ell}=\delta_{i\ell}$.
@@
Spickerman and Joyner [4] treated only the case $\mu=2$. We know that $A(2)=1$, hence $2$ is not a root of $A(x)$, whence
$\begin{split}c_{\ell}&=\frac{2^{k+1}-2^{k+1}+1}{(2-1)(2-\zeta_{\ell})}\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\frac{1}{\zeta_{\ell}^{k-1}}\\ &=\frac{1}{1/\zeta_{\ell}^{k}}\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\frac{1}{\zeta_{\ell}^{k-1}}\\ &=\zeta_{\ell}\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\end{split}$ (5.7)
Let us compare this to Spickerman and Joyner’s solution [4]. Because of the index shift by $1$, the relation of $w$ to $u$ is $w_{n}(\mu=2)=u_{n+1}$. Write
$u_{n}=\sum_{j=0}^{k-1}d_{\ell}\zeta_{k}^{n}\,.$ (5.8)
Our claim is $c_{\ell}\zeta_{\ell}^{n}=d_{\ell}\zeta_{\ell}^{n+1}$, or $c_{\ell}=d_{\ell}\zeta_{\ell}$. Note that Spickerman and Joyner wrote $\alpha_{j}$, whereas our notation is $\zeta_{\ell}$. From Spickerman and Joyner (unnumbered, second last line on p. 328 of [4])
$d_{\ell}=\zeta_{\ell}^{k}\frac{\zeta_{\ell}-1}{2\zeta_{\ell}^{k}-(k+1)}\,.$ (5.9)
Equate and compare
$\begin{split}\frac{c_{\ell}-d_{\ell}\zeta_{\ell}}{\zeta_{\ell}(\zeta_{\ell}-1)}&=\frac{1}{(k+1)\zeta_{\ell}-2k}-\frac{\zeta_{\ell}^{k}}{2\zeta_{\ell}^{k}-(k+1)}\\ &\propto(2\zeta_{\ell}^{k}-(k+1))-\zeta_{\ell}^{k}((k+1)\zeta_{\ell}-2k)\\ &=2\zeta_{\ell}^{k}-(k+1)-(k+1)\zeta_{\ell}^{k+1}+2k\zeta_{\ell}^{k}\\ &=2\zeta_{\ell}^{k}-2(k+1)\zeta_{\ell}^{k}+2k\zeta_{\ell}^{k}\\ &=0\,.\end{split}$ (5.10)
Our claim is validated.

### v4

@@ Abstract
We present new expressions for the $k$-generalized Fibonacci numbers, say $F_{k}(n)$. They satisfy the recurrence $F_{k}(n)=F_{k}(n-1)+\dots+F_{k}(n-k)$. Explicit expressions for the roots of the auxiliary (or characteristic) polynomial are presented, using Fuss-Catalan numbers. Properties of the roots are enumerated. We quantify the accuracy of asymptotic approximations for $F_{k}(n)$ for $n\gg 1$. Our results subsume and extend some results published by previous authors. We also present a basis (or ‘fundamental solutions’) to solve the above recurrence for arbitrary initial conditions. We comment on the use of generating functions and multinomial sums for the $k$-generalized Fibonacci numbers and related sequences. We note that the resulting multinomial sums are Dickson polynomials of the second kind in several variables. We also present what may be a new identity for companion matrices.
@@ 1 Introduction
The origin of this paper is somewhat curious. The author coauthored a paper [3] on the success run waiting times of Bernoulli trials. Recently, the author noticed that, for the special case $p=\frac{1}{2}$ ($p$ being the success parameter of the Bernoulli trials), elementary transformations yielded the solution for the so-called ‘$k$-generalized’ Fibonacci numbers, say $F_{k}(n)$. These numbers satisfy the recurrence
@@
The usual Fibonacci numbers are the special case $k=2$.
Our main contribution is to display explicit series solutions for all the roots of the characteristic equation of the above recurrence, for arbitrary $k\geq 2$ (leveraging formulas from [3]). Solutions in radicals are known for special cases such as the usual Fibonacci, tribonacci and tetranacci numbers. It has long been known that the characteristic equation has a unique positive real root (the ‘principal root’), whose magnitude exceeds that of all the other roots. Some authors, e.g. Wolfram [12], have derived expressions for the principal root and/or upper and lower bounds for it. We present expressions for all the roots and moreover enumerate several properties of the roots. For example, (i) there are no pure imaginary roots, and (ii) the smaller the real part of a root (less positive or more negative), the smaller its amplitude. We solve an associated Vandermonde system of equations and derive a Binet-style formula for the solution of the recurrence. (Other authors have also published the same or similar Binet-style formulas, e.g. Dresden and Du [4], but the roots had to be computed numerically, for example via a root-finding algorithm.)
We then apply our formalism to treat related problems published by other authors. For example, Spickerman and Joyner [11] solved the recurrence eq. (1.1), but with the initial values $1,1,2,4,8,\dots$. We generalize from powers of $2$ to powers of an arbitrary number $\mu$. Our solution is presented in Sec. 6.
Next, in Sec. 7 we present a set of ‘basis sequences’ or ‘fundamental solutions’ of the recurrence eq. (1.1). The usual scenario is to solve eq. (1.1) only for the initial conditions $F_{k}(0)=\dots=F_{k}(k-2)=0$ and $F_{k}(k-1)=1$, i.e. the initial $k$-tuple $(0,\dots,0,1)$. We present solutions for the other initial $k$-tuples $(1,0,\dots,0)$ through $(0,\dots,0,1,0)$. To the author’s knowledge, such a set of basis sequences has not been published in the literature. We leverage work from papers on solutions of linear recurrences, in particular by Wolfram [13].
Sec. 8 reviews the use of a companion matrix to solve linear recurrences with constant coefficients, mainly to make contact with the material in Sec. 9. In this context, employing a lemma by Wolfram [13], we present what may be a new identity for companion matrices. See Prop. 8.1.
Finally, in Sec. 9 we remark on the use of generating functions and multinomial sums to solve linear recurrences with constant coefficients. In this case, our presentation is more of an independent confirmation of known results, to demonstrate the application of our formalism for basis sequences. We treat the Narayana, Padovan and Perrin sequences. A referee kindly brought to the author’s attention that the multinomial sums in Sec. 9 are Dickson polynomials of the second kind in several variables.
2 Roots of the characteristic equation
Eq. (1.1) is a homogeneous linear difference equation and its characteristic equation is
@@ 2 Roots of the characteristic equation
The corresponding auxiliary (or characteristic) polynomial is $A(x)=x^{k}-x^{k-1}-\dots-1$. Several authors [12, 9, 10] have shown that the roots of $A(x)$ are simple. Let us denote the roots of $A(x)$ by $\zeta_{j}$, $j=0,\dots,k-1$. Then we express $F_{k}(n)$ as a weighted sum of powers of the roots as follows (i.e. a Binet-style formula)
$F_{k}(n)=\sum_{j=0}^{k-1}c_{j}\zeta_{j}^{n}\,.$ (2.2)
Here $\{c_{0},\dots,c_{k-1}\}$ is a set of coefficients which do not depend on $n$, but they do depend on $k$ and the initial conditions, i.e. the values of $F_{k}(0)$ through $F_{k}(k-1)$. It is standard to employ the initial conditions $F_{k}(0)=\dots=F_{k}(k-2)=0$ and $F_{k}(k-1)=1$. We shall begin with this, but later we shall treat alternative initial conditions, by Spickerman and Joyner [11]. In Sec. 7, we present a complete set of ‘fundamental solutions’ for the recurrence eq. (1.1), i.e. a basis which can be used to solve for arbitrary initial conditions.
Our main result is to present explicit expressions for the roots $\zeta_{j}$, in terms of so-called ‘Fuss-Catalan numbers’ (see the text by Graham, Knuth and Patashnik [6] for relevant definitions, formulas and identities). The connection is somewhat surprising. Dilworth and Mane [3] (hereafter ‘DM’) published a paper on the success run waiting times of Bernoulli trials. The recurrence for the probability mass function $f_{k}(n)$ is ([3] eq. (1)), with $p\in(0,1)$, q=1pq=1-p and $k\geq 2$,
@@
Set $p=q=\frac{1}{2}$ and $f_{k}(n)=F_{k}(n)/2^{n}$; then elementary manipulations yield eq. (1.1). Hence the solution for the $k$-generalized Fibonacci numbers $F_{k}(n)$ equals $2^{n}$ times the probability mass function $f_{k}(n)$ of success run waiting times of Bernoulli trials, with $p=q=\frac{1}{2}$. The auxiliary polynomial associated with eq. (2.3) is $\mathcal{A}_{p}(x)=x^{k}-qx^{k-1}-pqx^{k-2}-\cdots-p^{k-1}q$. Barry and Lo Bello [1] showed that the roots of $\mathcal{A}_{p}(x)$ are simple, which goes to show that there are fruitful connections between different fields of research. Let us denote the roots of $\mathcal{A}_{p}(x)$ by $\lambda_{j}$, $j=0,\dots,k-1$. Then $\zeta_{j}=2\lambda_{j}$, with $p=q=\frac{1}{2}$. The following properties of the roots were listed by DM ([3] Prop. 13).
  1. 1.
    It is known that $\mathcal{A}_{p}(x)$ has a unique positive root, which has a larger magnitude than all the other roots [5]. We denote it by $\lambda_{0}$ below.
  2. 2.
    If $k$ is odd, there are no other real roots. If $k$ is even, there is exactly one real negative root.
  3. 3.
    For all $k$, the remaining roots come in complex conjugate pairs.
  4. 4.
    There are no pure imaginary roots.
Wolfram [12] proved the first three results above for the case of the $k$-generalized Fibonacci numbers. The last item may not be so well-known. DM employed the terms ‘pricipal root’ for $\lambda_{0}$ and ‘secondary roots’ for the other $\lambda_{j}$, $j=1,\dots,k-1$.
@@
The principal root $\zeta_{0}$ is given by ([3] eq. (8))
@@
Wolfram’s solution is equivalent to our expression in eq. (2.7).
@@
It remains now to derive expressions for $F_{k}(n)$ for specific initial conditions, and also to derive asymptotic expressions for $n\gg 1$ (in which case the solution is dominated by the contribution from the principal root). To do so, it will be helpful below to work with the polynomial B(x)=(x1)A(x)B(x)=(x-1)A(x) for the following reason. Note that
@@ 3 Properties of roots
Distribution of the roots in the complex plane. Observe the following limits as $k\to\infty$: (i) for the principal root, $\limsup_{k\to\infty}\zeta_{0}=2$ (from eq. (3.3)) and (ii) for all the secondary roots, $\limsup_{k\to\infty}|\zeta_{j}|=1$ (from eq. (3.4)). Hence all the secondary roots approach the unit circle (from the inside), but the principal root approaches $2$ (from below). A graph of the roots in the complex plane is plotted in Fig. 1, for $k=40$. For $k\gg 1$, the secondary roots are also distributed approximately uniformly around the unit circle. A graph of the arguments of the secondary roots, i.e. $\arg(\zeta_{j})/(2\pi)$, is plotted in Fig. 2, for $k=40$. The points lie close to a straight line.
@@ 4 Vandermonde
First recall eq. (2.2). We set up a Vandermonde system of equations to solve for the coefficients $c_{j}$ as follows
[…]
We begin with the initial conditions $F_{k}(0)=\dots=F_{k}(k-2)=0$ and $F_{k}(k-1)=1$. (We treat the initial condition by Spickerman and Joyner [11] in Sec. 6.) The solution is ([3], eq. (60))
$c_{j}=\frac{1}{\prod_{s\neq j}(\zeta_{j}-\zeta_{s})}\,.$ (4.2)
@@
$B^{\prime}(x)=\frac{B(x)}{x-1}+\sum_{s=0}^{k-1}\frac{B(x)}{x-\zeta_{s}}\,.$ (4.4)
Then, noting that $B(\zeta_{j})=0$,
$B^{\prime}(\zeta_{j})=(\zeta_{j}-1)\prod_{s\neq j}(\zeta_{j}-\zeta_{s})\,.$ (4.5)
Also $B^{\prime}(x)=(k+1)x^{k}-2kx^{k-1}$, thus $B^{\prime}(\zeta_{j})=((k+1)\zeta_{j}-2k)\zeta_{j}^{k-1}$. Hence
$\begin{split}c_{j}&=\frac{1}{\prod_{s\neq j}(\zeta_{j}-\zeta_{s})}=\frac{\zeta_{j}-1}{B^{\prime}(\zeta_{j})}=\frac{\zeta_{j}-1}{((k+1)\zeta_{j}-2k)\zeta_{j}^{k-1}}\,.\end{split}$ (4.6)
@@
$F_{k}(n)=\sum_{j=0}^{k-1}\frac{\zeta_{j}-1}{(k+1)\zeta_{j}-2k}\zeta_{j}^{n-k+1}\,.$ (4.7)
@@
Remark 4.1.
The Binet-style sum in eq. (4.7) also works for $n<0$. The recurrence can be run backwards to negative values of $n$.
@@ 6 Alternative initial conditions: Spickerman and Joyner
Spickerman and Joyner [11] solved the recurrence eq. (1.1) as follows. First to avoid confusions of notation, we follow them and write $u_{n}$ in place of $F_{k}(n)$. They employed the initial conditions $(u_{0},\dots,u_{k-1})=(1,1,2,4,\dots,2^{k-2})$, i.e. $u_{0}=1$ and $u_{n}=2^{n-1}$ for $1\leq n\leq k-1$. Hence the next term is $u_{k}=1+1+2+4+\dots+2^{k-2}=2^{k-1}$. Let us change this slightly and shift the indices down by one (and discard $u_{0}$) and say instead $(v_{0},\dots,v_{k-1})=(1,2,4,\dots,2^{k-1})$, i.e. $v_{n}=2^{n}$ for $0\leq n\leq k-1$. Let us generalize this even more and replace $2$ by $\mu$, and write $w_{n}=\mu^{n}$ for $0\leq n\leq k-1$. Then the solution for $c_{j}$ in the Vandermonde matrix equations in Sec. 4 is easy; it is the Lagrange polynomial $c_{j}=L_{j}(\mu)$. The Lagrange polynomial $L_{j}(x)$ is
$L_{j}(x)=\frac{(x-\zeta_{0})\dots(x-\zeta_{j-1})(x-\zeta_{j+1})\dots(x-\zeta_{k-1})}{(\zeta_{j}-\zeta_{0})\dots(\zeta_{j}-\zeta_{j-1})(\zeta_{j}-\zeta_{j+1})\dots(\zeta_{j}-\zeta_{k-1})}\,.$ (6.1)
This can be obtained from the auxiliary polynomial, written in the form (see eq. (2.12)) $A(x)=(x^{k+1}-2x^{k}+1)/(x-1)$. The denominator of $L_{j}(x)$ is $D_{j}=A^{\prime}(\zeta_{j})$ and does not depend on $x$. The numerator of $L_{j}(\mu)$ is $N_{j}(\mu)=A(\mu)/(\mu-\zeta_{j})$, provided $\mu$ is not a root of $A(x)$. For now, we assume $\mu$ is not a root of $A(x)$. Then the numerator is
$N_{j}(\mu)=\frac{\mu^{k+1}-2\mu^{k}+1}{(\mu-1)(\mu-\zeta_{j})}\,.$ (6.2)
@@
The denominator is given by (note that the second term in $A^{\prime}(x)$ vanishes for $x=\zeta_{j}$)
$D_{j}=\frac{(k+1)\zeta_{j}^{k}-2k\zeta_{j}^{k-1}}{\zeta_{j}-1}=\frac{(k+1)\zeta_{j}-2k}{\zeta_{j}-1}\zeta_{j}^{k-1}\,.$ (6.4)
Hence the coefficient $c_{j}$ is
$c_{j}=\frac{N_{j}(\mu)}{D_{j}}=\frac{\mu^{k+1}-2\mu^{k}+1}{(\mu-1)(\mu-\zeta_{j})}\frac{\zeta_{j}-1}{(k+1)\zeta_{j}-2k}\frac{1}{\zeta_{j}^{k-1}}\,.$ (6.5)
@@
$\lim_{\mu\to 1}c_{j}=\frac{1-k}{(1-\zeta_{j})}\frac{\zeta_{j}-1}{(k+1)\zeta_{j}-2k}\frac{1}{\zeta_{j}^{k-1}}=\frac{k-1}{(k+1)\zeta_{j}-2k}\frac{1}{\zeta_{j}^{k-1}}\,.$ (6.6)
@@
What if $\mu$ is a root of $A(x)$? Actually this is easy. If $\mu=\zeta_{i}$ then by construction of the Lagrange polynomials, $L_{i}(\zeta_{i})=1$ and $L_{j}(\zeta_{i})=0$ for $i\neq j$. Hence $c_{i}=1$ and $c_{j}=0$ for $i\neq j$, i.e. $c_{j}=\delta_{ij}$.
Proposition 6.1.
Our expression in eq. (6.5), evaluated for $\mu=2$, is equivalent to Spickerman and Joyner’s solution [11].
Proof.
@@
Spickerman and Joyner treated only the case $\mu=2$. We know that $A(2)=1$, hence $2$ is not a root of $A(x)$, whence
$\begin{split}c_{j}&=\frac{2^{k+1}-2^{k+1}+1}{(2-1)(2-\zeta_{j})}\frac{\zeta_{j}-1}{(k+1)\zeta_{j}-2k}\frac{1}{\zeta_{j}^{k-1}}\\ &=\frac{1}{1/\zeta_{j}^{k}}\frac{\zeta_{j}-1}{(k+1)\zeta_{j}-2k}\frac{1}{\zeta_{j}^{k-1}}\\ &=\zeta_{j}\frac{\zeta_{j}-1}{(k+1)\zeta_{j}-2k}\,.\end{split}$ (6.7)
The expression for $u_{n}$ by Spickerman and Joyner is (unnumbered, second last line on p. 328 of [11])
$u_{n}=\sum_{j=0}^{k-1}\frac{(\alpha_{j}^{k+1}-\alpha_{j}^{k})\alpha_{j}^{n}}{2\alpha_{j}^{k}-(k+1)}\,.$ (6.8)
They denoted the roots by $\alpha_{j}$, whereas we say $\zeta_{j}$. We thus express eq. (6.8) in the form $u_{n}=\sum_{j=0}^{k-1}d_{j}\zeta_{j}^{n}$, where
$\begin{split}d_{j}&=\frac{(\zeta_{j}^{k+1}-\zeta_{j}^{k})}{2\zeta_{j}^{k}-(k+1)}\\ &=\zeta_{j}^{k}\frac{\zeta_{j}-1}{2\zeta_{j}^{k}-(k+1)}\,.\end{split}$ (6.9)
Because of the index shift of $n$ by $1$, the relation of $w$ to $u$ is $w_{n}(\mu=2)=u_{n+1}$. Our claim is therefore $c_{j}\zeta_{j}^{n}=d_{j}\zeta_{j}^{n+1}$, or $c_{j}=d_{j}\zeta_{j}$. Employing eqs. (6.7) and (6.9) yields
$\begin{split}\frac{c_{j}-d_{j}\zeta_{j}}{\zeta_{j}(\zeta_{j}-1)}&=\frac{1}{(k+1)\zeta_{j}-2k}-\frac{\zeta_{j}^{k}}{2\zeta_{j}^{k}-(k+1)}\\ &\propto(2\zeta_{j}^{k}-(k+1))-\zeta_{j}^{k}((k+1)\zeta_{j}-2k)\\ &=2\zeta_{j}^{k}-(k+1)-(k+1)\zeta_{j}^{k+1}+2k\zeta_{j}^{k}\\ &=2\zeta_{j}^{k}-2(k+1)\zeta_{j}^{k}+2k\zeta_{j}^{k}\\ &=0\,.\end{split}$ (6.10)
Our claim is validated. ∎